### v1

# Calibrating Translation Decoding with Quality Estimation on LLMs

Abstract
Neural machine translation (NMT) systems typically employ maximum a posteriori (MAP) decoding to select the highest-scoring translation from the distribution mass. However, recent evidence highlights the inadequacy of MAP decoding, often resulting in low-quality or even pathological hypotheses — the decoding objective is not aligned with real-world translation quality. This paper proposes calibrating hypothesis likelihood with translation quality from a distributional view by directly optimizing their Pearson correlation — thereby enhancing the effectiveness of translation decoding. With our method, translation on large language models (LLMs) improves substantially after limited training (2K instances per direction). This improvement is orthogonal to those achieved through supervised fine-tuning, leading to substantial gains across a broad range of metrics and human evaluations — even when applied to top-performing translation-specialized LLMs fine-tuned on high-quality translation data, such as Tower, or when compared to recent preference optimization methods, like CPO. Moreover, the calibrated translation likelihood can directly serve as a strong proxy for translation quality, closely approximating or even surpassing some state-of-the-art translation quality estimation models, like CometKiwi. Lastly, our in-depth analysis demonstrates that calibration enhances the effectiveness of MAP decoding, thereby enabling greater efficiency in real-world deployment. The resulting state-of-the-art translation model, which covers 10 languages, along with the accompanying code and human evaluation data, has been released to the community: https://github.com/moore3930/calibrating-llm-mt.
1 Introduction
The training of Neural Machine Translation (NMT) has long been formulated as a maximum likelihood estimation (MLE) problem, and maximum a posteriori decoding (MAP) is a common decision rule, aiming to identify the highest-scoring translation. However, recent findings in NMT suggest that such a system has serious flaws. One particularly counterintuitive issue is the beam search curse (Koehn and Knowles, 2017; Murray and Chiang, 2018; Ott et al., 2018; Kumar and Sarawagi, 2019), where translation quality deteriorates as search approximations improve — higher-probability hypotheses are possibly worse translations.
Ott et al. (2018) explain this as a calibration issue — there is a generally low correlation between hypothesis likelihood and quality beyond a certain likelihood value. The flaws of poorly calibrated translation models can be categorized into two main aspects. First, their performance tends to be suboptimal, as better hypotheses may remain hidden within the distribution mass, inaccessible to MAP decoding (Ott et al., 2018). Second, likelihood, as a key measure of uncertainty, can serve as a valuable indicator of translation errors within the output at test time (Wang et al., 2020; Fomicheva et al., 2020); however, poorly calibrated models cannot effectively support this use.
Although these investigations were conducted on encoder-decoder architectures, we found that they also apply to translations elicited from LLMs. More specifically, we observed that (1) LLMs generally exhibit poor likelihood-quality calibration in translation, even when scaled up in capacity or continued pretrained with large amounts of multilingual data to enhance multilinguality; and (2) despite instruction fine-tuning with high-quality data leads to clear improvements in translation quality, as recommended by current best practices (Xu et al., 2023; Rei et al., 2024a), improvements in calibration are limited and sometimes barely observable.
Prior studies have tried to mitigate this miscalibration issue by introducing an additional optimization step during inference time, as shown in Figure 1 upside, known as quality-aware decoding (QAD) (Fernandes et al., 2022). These approaches typically involve generating multiple candidate translations through sampling, followed by reranking or voting using reference-free and/or reference-based machine translation metrics, such as Best-of-N (BoN) sampling (Rei et al., 2024a; Faria et al., 2024; Brown et al., 2024; Ichihara et al., 2025) and Minimum Bayes Risk (MBR) decoding (Kumar and Byrne, 2004; Freitag et al., 2022a). While effective, this line of methods inevitably incurs significantly higher latency. For example, the current top-performing system in the WMT24 competition (Kocmi et al., 2024a), Tower (Rei et al., 2024a), requires sampling 100 times for each prompt from a 70B LLM-based translation model; this is impractical, particularly for online environments.
This paper proposes a simple yet highly effective method to directly optimize the calibration during training, where we encourage the likelihoods of translation hypotheses to align more closely with their quality distribution by directly optimizing Pearson correlations between them. More specifically, as shown in Figure 1 lower side, given a translation prompt $x$, we sample multiple hypotheses from an NMT model during training time and compute both their likelihoods and quality scores, as measured by any external metric, such as COMET. We then define the loss as the negative Pearson correlation between the two sets of data points and minimize it directly using a standard gradient-based optimizer. While simple, our approach yields substantial performance improvements with limited training — using only 2K instances per language — even when applied to state-of-the-art LLM-based MT systems such as Tower. Moreover, due to clear calibration improvements, the resulting models’ likelihood can directly serve as a strong proxy for translation quality, even surpassing some state-of-the-art translation quality estimation (QE) models, like Comet-Kiwi (Rei et al., 2022a).
Figure 1: The illustration of quality-aware decoding (upper side) and our calibration method (lower side). The former explores the decoding space by adding an extra step during test-time decoding, which involves multiple rounds of sampling followed by reranking or voting to select the highest-scoring output. Our method focuses on training time optimization, intending to calibrate the distribution of hypothesis likelihoods to the corresponding qualities.
Our main contributions are summarized as follows:
  • We propose a simple yet effective method to mitigate the well-known miscalibration issue in translation during training time, where we calibrate translation decoding for quality from a holistic view by directly optimizing the Pearson correlation objective. With limited training, our models consistently outperform the strongest LLM-based MT systems – such as ALMA (Xu et al., 2023, 2024) and Tower (Rei et al., 2024a; Alves et al., 2024) series as well as the GPT-4o model — by a substantial margin across multiple automatic and human evaluation metrics. Extensive experiments demonstrate effectiveness over recent preference optimization methods, such as CPO (Xu et al., 2024), for translation.
  • Our method, to some extent, unifies quality optimization and quality estimation (QE) in translation by sharing one single objective. Specifically, the uncertainty quantification of our model’s output, measured by average log-likelihood, exhibits a strong correlation with human judgments, which even outperforms some state-of-the-art QE models in the community that were explicitly trained with human-annotated data, such as CometKiwi (Rei et al., 2022b). This finding supports the view that a well-performing model should inherently “know” what constitutes a good translation.
  • Our in-depth analysis demonstrates that improved likelihood-quality calibration — achieved through our calibration method — enhances the effectiveness of maximum a posteriori decoding, offering greater efficiency potential for real-world deployment.
2 Background
At a practical level, this study aims to: (1) optimize translation quality, and (2) calibrate the model’s uncertainty quantification, i.e., likelihood, with human judgments. To this end, we briefly introduce key foundational concepts, including metric meta-evaluation and uncertainty in translation.
Translation metric meta-evaluation. To automatically assess translation quality, various metrics have been proposed over the past decades, such as BLEU (Papineni et al., 2002), ChrF (Popović, 2015), COMET (Rei et al., 2020, 2022b), MetricX (Juraska et al., 2023, 2024). Metric meta-evaluation assesses how well these metrics correlate with human judgments, where a few well-known statistics are applied, such as Pearson’s $r$, Spearman’s $\rho$, and Kendall’s $\tau$. In this paper, we focus on calibrating hypothesis likelihood with translation quality by directly optimizing Pearson correlation, enabling simultaneous quality optimization and estimation within a single model; see §3. To ensure robust cross-metric evaluation, we report Spearman’s and Kendall’s scores for quality estimation; see §5.2.
Translation uncertainty. Uncertainty estimation (Ott et al., 2018; Huang et al., 2024) is an important diagnostic tool often used to assess confidence in the model’s output. Previous work (Galley et al., 2015; Fomicheva et al., 2020) has shown that, under a Bayesian framework, appropriate quantification, such as entropies, posterior distributions, or attention weights, can yield uncertainty estimates that correlate well with human judgments of translation quality, even without human-labeled training data. In this paper, we focus on a specific form of quantification — likelihood — due to its natural connection to translation training and decoding.
Previous approaches.
A few different themes investigate optimizing translation performance using signals from downstream quality evaluation metrics. We briefly summarize as follows.
  1. 1.
    Test-time quality optimization. To alleviate the gap between decoding objective and translation quality, Minimum Bayes Risk (MBR) decoding (Kumar and Byrne, 2004; Freitag et al., 2022a; Yang et al., 2024) suggests basing decoding decisions on statistics gathered from the distribution holistically. As reference-free quality estimation (QE) progresses, Best-of-N sampling (Rei et al., 2024b; Faria et al., 2024; Brown et al., 2024; Ichihara et al., 2025) with a QE model becomes a straightforward strategy. However, both of these methods will result in a significant amount of additional decoding time. Moreover, the risk of metric bias increases, as translation improvements may stem from metric hacking rather than genuine quality enhancements (Skalse et al., 2022; Kovacs et al., 2024).
  2. 2.
    Preference optimization. Recently, direct preference optimization (DPO) and its variants (Rafailov et al., 2023; Xu et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024) have emerged as promising approaches for post-training LLMs, providing efficient alternatives to reinforcement learning from human feedback (RLHF). This line of methods often employs pairwise preferences under the Bradley-Terry framework (Bradley and Terry, 1952) to provide rewards, or a more general Plackett-Luce ranking framework (Plackett, 1975) when multiple ranked hypotheses are available — a few studies (Xu et al., 2024; Zhu et al., 2024) in translation also follow them. Additionally, some works (Guo et al., 2024; Lambert et al., 2024) employ these methods in an on-policy training paradigm for online LLMs alignments. Notably, regardless of employing pairwise/listwise data or on-/off-policy training, the objective is generally to maximize expected rewards. One key distinction of our method, as will be shown in §3, is that our training objective is to maximize correlation instead of expected rewards. Therefore, we follow prior research (Ott et al., 2018; Zhao et al., 2022) and use the term "calibration" in this paper to highlight the difference. More detailed discrepancies can be found in the next section.
3 Translation Calibration
Formally, given a parameterized auto-regressive language model $\theta$, and a translation instruction $x$, the log-likelihood of each translation hypothesis $y_{i}$ is denoted as $p_{\theta}(y_{i}|x)$. Meanwhile, the quality of this translation can be defined as $q(y_{i}|x)$ where $q$ represents any external quality evaluation model. When sampling hypotheses $y_{i}$ from $\theta$ conditioned on $x$, both $p_{\theta}(\cdot|x)$ and $q(\cdot|x)$ can be viewed as random variables defined over the output space.
Our goal is to calibrate the likelihoods of generated hypotheses with their quality — formally, to maximize the correlation between $p_{\theta}(\cdot|x)$ and $q(\cdot|x)$. Here, we bring the statistic, i.e., Pearson score $\rho(P,Q)$, to measure it. Pearson score $\rho(P,Q)$ measures the linear correlation between two variables $P$ and $Q$, which is defined as follows:
[…]
where $\mathrm{cov}(P,Q)$ is the covariance between $\mathrm{P}$ and $\mathrm{Q}$, while $\sigma_{P}$ and $\sigma_{Q}$ are the corresponding standard deviations, and $\mu_{p}$ and $\mu_{q}$ are the corresponding means. This formulation computes the correlation by integrating the product of the centered values of $p$ and $q$ normalized by their standard deviations. Due to its scale-invariant properties and ability to measure trend consistency, the Pearson correlation coefficient is widely used in areas such as translation metric meta-evaluation.
Given the intractably large decoding space, we employ sampling to approximate it. For each source sentence $x$, we first generate $k$ hypotheses $y_{i}$ ($i\in\{1,\dots,k\}$) by repeatedly prompting a large language model $\theta$ with nuclear sampling, and compute the corresponding $p_{\theta}(y_{i}|x)$. Also, we use a third-party metric model, like CometKiwi, to evaluate each hypothesis’s quality, i.e., $q(y_{i}|x)$.
By independently normalizing $p_{\theta}(\cdot|x)$ and $q(\cdot|x)$ to have zero mean and unit variance, we can define Pearson-based training loss for each source sentence $x$ as follows:
$\mathcal{L}_{\text{pearson}}=-\frac{1}{k}\sum_{\begin{subarray}{c}i=1\\ y_{i}\sim p_{\theta}(y\mid x)\end{subarray}}^{k}\left(\frac{p_{\theta}(y_{i}\mid x)-\mu_{p}}{\sigma_{p}}\cdot\frac{q(y_{i}\mid x)-\mu_{q}}{\sigma_{q}}\right).$ (2)
Note that we also introduce a supervised fine-tuning (SFT) term on the highest-scoring samples as a regularizer to ensure that the model’s likelihood distribution remains grounded in high-quality translations, as the Pearson objective alone only enforces correlation and does not constrain the absolute scale. The final loss for calibration is formulated as $\mathcal{L}_{\text{cal}}=\mathcal{L}_{\text{pearson}}+\mathcal{L}_{\text{sft}}$.
An off-policy formulation can be obtained by trivially replacing the current model $\theta$ with an external model $\theta^{\ast}$ for sampling. Overall, by minimizing $\mathcal{L}_{\text{cal}}$ here, we encourage the Pearson score between $p$ and $q$ to be higher. In practice, we use a gradient-based optimizer, Adam, to optimize $\theta$ for this goal. Despite its simplicity, several important characteristics are captured in this formulation:
  • It models hypothesis qualities from a holistic view, enabling the model to make finer-grained distinctions in translation quality within decoding space.
  • It considers the value of translation quality by the metric function $q(\cdot|x)$, which is ignored in virtually all existing methods based on Bradley-Terry and Plackett-Luce, such as CPO.
  • Pearson’s correlation inherently applies normalization on a group of both likelihood and quality points. This normalization makes the objective invariant to scale and shift, thereby promoting stable and robust optimization across diverse input distributions.
  • The objective — Pearson’s score itself – is inherently shared with that of translation metric meta-evaluation, offering a unified perspective for both quality optimization and estimation.
Overall, the Pearson loss can be reduced to a mere dot product between two sets of normalized points. Despite this extreme simplicity, we show in the next section that it is highly effective in simultaneously optimizing translation quality and quality estimation within a single model.
4 Experimental Settings
4.1 Base Models
We base our experiments on two series of strong translation-specified LLMs, i.e., ALMA-Base (Xu et al., 2023) and Tower-Base (Alves et al., 2024), in both 7B and 13B variants. Both models are obtained by continued pretraining of LLAMA2-7B and LLAMA2-13B on extensive multilingual data. A key distinction is that ALMA-Base relies exclusively on monolingual data (Xu et al., 2023) for continued pretraining, whereas Tower-Base incorporates both parallel and monolingual data to improve cross-lingual capabilities. Both models achieve remarkable translation performance. For instance, the Tower series systems (Rei et al., 2024b) won all of the translation tasks in the WMT24 general translation tracks (Kocmi et al., 2024a), showing state-of-the-art performance.
4.2 Data and Evaluation
Translation evaluation dataset. For fair comparison with the ALMA and Tower model series, we evaluate translation performance on the WMT22 and WMT24 datasets (Zerva et al., 2022; Kocmi et al., 2024a), respectively. Detailed dataset descriptions can be found in Appendix A. Following current best practice (Freitag et al., 2022b, 2023), all of the results in this paper are measured by current widely used neural metrics, involving reference-based metric COMET, and the reference-free metrics XCOMET, CometKiwi-XL, and CometKiwi-XXL [^1] The corresponding metric model versions are Unbabel/wmt22-comet-da, Unbabel/XCOMET-XXL, Unbabel/wmt23-cometkiwi-da-xl, and Unbabel/wmt23-cometkiwi-da-xxl, respectively..
Metric meta-evaluation dataset. To assess how well translation likelihoods correlate with human judgments, we use the sentence-level quality estimation (QE) datasets during WMT-22 (Zerva et al., 2022), where sentence pairs are annotated in multiple ways to reflect translation quality, e.g., direct assessments (DA), post-edits (PE), and multidimensional quality metrics (MQM) (Freitag et al., 2021). In this paper, we use MQM scores as the ground-truth human annotations against which the metrics’ scores are evaluated, as MQM scores come from expert translators and are more reliable than the crowdsourced DA scores. Detailed description for the dataset can be found in Appendix A.
Calibration dataset. For the training set, we merge all English sentences from the Flores-200 dataset (Costa-Jussà et al., 2022) in dev and devtest splits and use them as the source, consisting of 2009 samples. For both on- and off-policy experiments, we use these sentences to construct translation prompts for each direction. The prompt templates can be found in Appendix E. For off-policy setting, we query gpt-4o-mini 16 times per prompt, employing nucleus sampling with a temperature of 1.0 and a top-p of 0.98. The resulting bitexts are evaluated using varying metrics to reflect corresponding quality scores. For on-policy experiments, all settings remain the same, except that a top-k value of 5 is used for each sampling step, and the model itself is queried directly.
4.3 Training Setups
For all experiments, we train the models using LoRA (Hu et al., 2022) with a rank of 8. We set $\alpha$ to 32 and use a dropout rate of 0.05. The training process involves a batch size of 32, a gradient accumulation step of 8, and a maximum length of 512 tokens for each sequence. We train each model for 3 epochs and select checkpoints based on the best performance measured by XCOMET on the validation set, NTREX (Federmann et al., 2022), which consists of 1997 samples for each direction. For more reliable experiments, we train models with learning rates ranging from 1e-5 to 1e-4, and report the results using the best-performing learning rate for each setting. We use Adam (Kingma and Ba, 2014) as the optimizer. Unless otherwise specified (e.g., § 2), we use CometKiwi-XXL as the signal during training and report results in all other metrics such as XCOMET, COMET, and CometKiwi-XL. All experiments are conducted using H100 GPUs, with 7B models trained on a single GPU and 13B models trained on two GPUs.
5 Results
5.1 Calibration Leads to Clear Quality Improvements
Table 1: Evaluation of enxx translation on WMT24 using CometKiwi-XL and XCOMET. Results are reported for all languages covered during Tower-v1 pretraining. Note that the Tower-v2 models, including Tower-70B-v2, have not been publicly released. We report their best results as published by Rei et al. (2024a). For GPT-4o and GPT-4o-mini, we use the prompts following  Hendy et al. (2023). Results in other metrics can be found in Appendix B.1. Notably, according to Kocmi et al. (2024b), improvements of $\geq 1.99$ in XCOMET or $\geq 0.94$ in COMET scores correspond to at least 90% estimated accuracy in human judgment — both of which are achieved by our method.
ende enes enru enzh enfr
Models KIWI-XL XCOMET KIWI-XL XCOMET KIWI-XL XCOMET KIWI-XL XCOMET KIWI-XL XCOMET
Closed GPT-4o-mini 68.3 91.7 70.2 87.0 68.1 81.6 69.0 79.7 65.6 83.0
GPT-4o 68.6 92.6 70.6 87.7 69.1 83.4 69.9 81.3 66.0 83.9
Tower-70B-v2
Tower-70B-v2 + MBR/TRR 72.3 74.5 74.2 72.6
TowerInstruct-7B 69.0 91.7 70.8 86.9 69.0 81.5 68.5 78.7 67.9 84.1
TowerBase-7B
     + SFT on BoN data 70.0 92.0 70.8 86.5 69.6 81.6 68.4 77.9 68.0 83.7
     + CPO 71.1 93.1 72.0 87.6 71.6 83.8 70.4 80.9 69.3 85.8
     + Calibration (ours) 71.6 93.6 73.5 89.0 72.4 84.8 70.4 81.0 70.0 86.8
[…]
TowerBase-13B
     + SFT on BoN data 71.1 92.7 71.8 87.5 71.3 82.8 70.1 80.0 68.0 84.4
     + CPO 70.5 92.2 72.0 87.7 71.9 84.0 70.3 81.4 68.8 85.5
     + Calibration (ours) 72.5 94.2 73.8 90.0 73.6 86.4 72.1 83.6 70.8 87.5
[…]
     + SFT on BoN data 70.7 92.7 71.8 87.1 70.8 82.9 70.5 80.4 68.5 84.4
     + CPO 71.2 93.0 73.1 89.0 72.3 85.1 71.8 83.6 70.0 86.9
     + Calibration (ours) 72.4 94.0 73.9 89.9 73.6 86.1 72.6 83.7 70.8 87.4
ennl enit enpt enko Avg.
Models KIWI-XL XCOMET KIWI-XL XCOMET KIWI-XL XCOMET KIWI-XL XCOMET KIWI-XL XCOMET
Closed GPT-4o-mini 69.4 88.9 68.1 83.7 71.2 87.6 73.2 84.2 69.2 85.3
GPT-4o 70.6 90.5 68.7 85.7 71.5 88.5 73.7 85.6 69.8 86.6
Tower-70B-v2
Tower-70B-v2 + MBR/TRR
TowerInstruct-7B 71.5 90.9 71.1 86.1 71.1 86.8 73.6 82.8 70.3 85.5
TowerBase-7B
     + SFT on BoN data 71.5 89.6 70.8 85.4 72.5 87.6 75.7 84.1 70.8 85.4
     + CPO 71.9 90.9 72.2 86.7 73.4 88.7 76.1 87.2 72.0 87.2
     + Calibration (ours) 73.3 91.9 73.5 88.1 74.8 89.9 76.8 87.2 72.9 88.0
[…]
TowerBase-13B
     + SFT on BoN data 71.7 90.4 71.6 86.1 73.0 88.1 76.2 85.2 71.6 86.4
     + CPO 72.3 90.8 72.5 87.4 72.2 86.9 76.9 87.9 71.9 87.1
     + Calibration (ours) 73.9 92.6 73.9 89.3 75.2 90.4 78.0 89.5 73.8 89.3
[…]
     + SFT on BoN data 72.3 90.7 71.6 86.2 72.7 87.9 76.2 86.0 71.7 86.5
     + CPO 73.3 92.3 73.1 88.5 74.0 89.7 77.4 89.3 72.9 88.6
     + Calibration (ours) 74.2 93.2 74.1 89.6 75.1 90.7 78.1 89.7 73.9 89.4
In this section, we demonstrate the effectiveness of our calibration approach by applying it to some state-of-the-art translation systems, e.g., Tower and ALMA series. Table 1 presents the results for Tower series under an off-policy setting (see §4.2 for the training data), measured by CometKiwi-XL, the official reference-free metric for WMT24, and XCOMET, the best performing metric as evaluated by  Freitag et al. (2023). TowerInstruct, +SFT, and +Calibration in the table denote the results decoded by beam search with a beam size of 5. TowerInstruct-7B/-13B, and TowerInstruct-Mistral-7B are official implementations (Rei et al., 2024a), supervised fine-tuned on the corresponding base models using TowerBlock, a set of high-quality translation instructions constructed through careful data selection and filtering from human-annotated translation datasets, consisting of 637k instructions.
We also conducted supervised fine-tuning on the TowerBase series using 2K Best-of-N samples per direction, selected from our calibration dataset (see §4.2) based on the highest CometKiwi-XXL scores. It is easy to see that the resulting average performances are comparable to those of official Instruction models, highlighting the quality of our dataset. Note that performance degrades when fine-tuning on the full calibration set, as some of the sampled bitext examples are of low quality.
When applying our calibration approach, very strong improvements can be observed across all directions, metrics, and base models, for example: (1) It leads to an average improvement of +2.8 points in KIWI-XL and +2.7 points in XCOMET over TowerInstruct-Mistral-7B. Additionally, Table 3 shows gains of +3.6 points in KIWI-XXL and +1.2 points in COMET, respectively; and (2) This performance is comparable to that of the current top-performing system, that is, Tower-70B-v2 equipped with 100-time-sampling MBR/TRR [^2] TRR (Rei et al., 2024a) denotes an ensemble strategy that applies reranking based on multiple metric model to select the best candidate from 100 sampled hypotheses. The authors report TRR results when it surpass MBR. — while being approximately 200× speed up [^3] We roughly estimate the latency of the Tower-70B-v2 model to be 10 times that of the Tower-Mistral-7B model. Meanwhile, the former employs 100× sampling, while the latter uses beam search with a beam size of 5..
We also compare with CPO (Xu et al., 2024), a widely used preference optimization method in translation, following their original setting and selecting the highest- and lowest-scoring sampled candidates as accepted and rejected samples, respectively — consistent and substantial improvements are observed. Additionally, Appendix B.2 reports results based on the ALMA base model, showing gains of +2.0 XCOMET and +1.4 KIWI-XXL for out-of-English translation over CPO.
Note that experimental results in on-policy settings are also provided in Appendix B.3, where substantial improvements can also be consistently observed across metrics, demonstrating effectiveness in different training dynamics. Unless otherwise specified, the following analysis focuses on the off-policy setting to simplify the investigation of the properties of our calibration method.
5.2 Calibration Leads to Direct Quality Estimation
As detailed in §3, we shared the objective for translation quality optimization and estimation, although supervisions are from machine annotations instead of human annotations. If optimized effectively, the resulting model should inherently acquire the ability to assess translation quality using hypothesis log-likelihood as a metric. This section evaluates how effectively calibration can elicit this capability.
We use the WMT22 metric meta-evaluation dataset (Zerva et al., 2022) and follow the official practice, using the multidimensional quality metrics (MQM) scores as golden quality annotations, which are provided by expert translators and are considered more reliable than other annotations. We evaluate all training directions on Tower that overlap with the WMT dataset, namely, ende and enru.
[…]
Figure 2 depicts the Spearman score (metric performance) and the corresponding translation performance under different settings for Tower-7B and Tower-13B, including: (1) supervised fine-tuning using varying amounts of best-of-N samples (400/800/1200/1600/2000 samples per direction), (2) scaling the base model size from 7B to 13B, and (3) applying our calibration method. It shows that:
  • The TowerInstruct-7B/-13B models exhibit relatively low performance in both the translation and quality estimation abilities. The resulting models manifest progressively improved translation performance as more best-of-N samples are included for supervised fine-tuning. Interestingly, the corresponding quality estimation abilities, measured by Spearman scores calculated on the WMT22 dataset, increase from around 51.5 to 54.0 points. We attribute this to fine-tuning on Best-of-N samples, which are predominantly high-quality, encouraging the model to assign higher likelihoods to better hypotheses.
  • Examining the effects of scaling, we observe that: (1) scaling up from 7B to 13B generally improves translation performance for both the original TowerInstruct models and the fine-tuned models using our best-of-N data; (2) however, its impact on calibration — that is, the models’ quality estimation ability — remains minimal.
  • Our calibration method manifests very strong improvements in both translation and quality estimation. For example, when applying our method to TowerBase-13B, the resulting model surpasses some state-of-the-art systems in both translation performance and quality estimation ability, i.e., Tower-70B-v2+MBR/TRR and CometKiwi, simultaneously.
Results in other statistics are provided in Appendix C. Overall, we observe a clear, albeit sometimes non-linear, correlation between the models’ translation performance and their quality estimation ability. These results suggest, to some extent, a unified perspective: a well-performing translation system should inherently "know" what constitutes a good translation. In turn, we also suggest optimizing translation quality by improving calibration on LLMs, rather than relying solely on extreme scaling or supervised fine-tuning, as the latter approaches show relatively limited effectiveness.
5.3 Calibration Enhances Effectiveness for Maximum a Posteriori Decoding
Figure 3: The average performance (ende,es,ru,zh) measured by COMET score for both TowerInstruct-Mistral-7B and our calibrated models, when (1) varying sampling size for Best-of-N sampling and (2) applying beam search with a beam size of 5.
Compared to test-time optimization, such as Best-of-N sampling, our approach performs sampling and directly calibrates the likelihoods of hypotheses with their translation qualities during training. Ideally, for a well-calibrated model, the potential of maximum a posteriori (MAP) decoding — such as beam search — should be more effectively realized.
Figure 3 compares the effectiveness of beam search (with a beam size of 5) to that of Best-of-N sampling with varying sampling sizes, for both the baseline (TowerInstruct-Mistral-7B) and our calibrated models. To ensure fairness, we adopt a cross-metric evaluation for Best-of-N sampling: candidates are reranked using KIWI-XXL, while best-scoring results are evaluated using COMET. Two key findings are clearly illustrated in Figure 3: (1) As the Best-of-N sampling size increases, translation performance improves for both models, with our model consistently outperforming the baseline across all settings. (2) Under the MAP decoding strategy (i.e., beam search), the baseline model (dashed blue line) underperforms compared to Best-of-N sampling with a sampling size as small as 10. However, after calibration, beam search achieves performance (dashed red line) comparable to that of Best-of-N sampling with a size of 100, showing much stronger realizations of MAP capabilities. Note that beam search is dozens of times more efficient than Best-of-N sampling.
Overall, compared to the state-of-the-art translation system, our calibration method yields clear performance improvements and offers greater efficiency potential for real-world deployment. Unlike computation-intensive test-time optimization methods like Best-of-N sampling, it inherently improves effectiveness by calibrating decoding to the real-world quality objective, unlocking the potential of efficient MAP decoding rules, such as beam search.
6 Analysis
6.1 Cross-Metric Evaluation
Table 2: Average score differences on WMT24 over those of TowerInstruct-Mistral-7B across metrics, when applying (a) supervised fine-tuning and (b) our calibration method under different settings.
(a) SFT Over TowerInstruct-Mistral-7B
Objective KIWI-XXL XCOMET COMET KIWI-XXL +0.3 -0.2 -0.3 XCOMET -0.3 +0.1 -0.3 COMET -0.5 -0.7 -0.4
(b) Calibration Over TowerInstruct-Mistral-7B
Objective KIWI-XXL XCOMET COMET KIWI-XXL +3.5 +2.7 +1.2 XCOMET +2.4 +2.8 +1.0 COMET +2.1 +1.4 +0.8
A potential concern with directly optimizing towards translation quality is metric gaming (Casper et al., 2023; Kovacs et al., 2024). To address this, we conduct experiments with TowerInstruct-Mistral-7B, using various metrics to assess hypothesis quality during training and to examine whether improvements are consistent across metrics. To maximize the contrast between metrics, we focus on three representative ones in our setting: (1) COMET, a strong widely used reference-based metric; (2) KIWI-XXL, the strongest reference-free metric in the KIWI series; and (3) The reference-free version of XCOMET that combines sentence-level evaluation with error span detection.
Table 2 presents the average improvements on WMT24 when applying (a) supervised fine-tuning and (b) our calibration method, using these three different metrics. It shows that: (1) When fine-tuning on our Best-of-N dataset, we generally observe slight performance degradation for all settings, except when KIWI-XXL and XCOMET are used simultaneously as both the training objective and the evaluation metric. We attribute the degradation to the fact that TowerInstruct-Mistral-7B is already stronger than GPT-4o-mini, the system we queried. Meanwhile, we attribute the score improvements in KIWI-XXL to the existence of slight metric gaming. (2) When applying our calibration methods, we observe consistently strong improvements across all training objectives and evaluation metrics, highlighting cross-metric effectiveness. (3) The reference-based metric, COMET, demonstrates relatively low effectiveness during training in both settings. Conservatively, we attribute this to the fact that reference-based metrics constrain the feasible space of positive hypotheses to be similar to the reference, thus weakening the expressiveness of the quality space.
6.2 Human Study
Human Evaluation Summary
TScore CScore Win Loss Tie
4.77 5.49 106 32 62
Figure 4: Human evaluation results in enzh direction: score distribution and an aggregated summary. TScore and CScore denote the average scores over 200 annotations for TowerInstruct-Mistral-7B and our calibrated model, respectively. Win, Loss, and Tie indicate the number of samples for which our model outperformed, underperformed, or tied with TowerInstruct-Mistral-7B.
The preceding results demonstrate the effectiveness of our method across different settings and metrics. Here, we incorporate human evaluation as direct evidence for comprehensive improvements.
We randomly sample 200 translation outputs in the directions of enzh/de/ru for both the baseline model (TowerInstruct-Mistral-7B) and our calibrated model from the WMT24 dataset. Several bilingual speakers, native in the respective target languages, were recruited to annotate each translation on a 0–6 scale, following the criteria outlined in (Kocmi et al., 2022), see Appendix D.1. Annotation was performed using integer scores rather than continuous real values.
Figure 4 shows the overview results for enzh annotations, covering both fine-grained score distributions and an aggregated summary. It shows that our calibrated model yields significantly more near-perfect translations (137 vs. 67 for scores of 6), with the main gains over the baseline observed in improved handling of minor-to-major errors (denoted as score of 3/4/5) in grammar, wording, or tone, see Appendix D.2. Overall, the average score for the calibrated model is 5.49 out of 6, showing very strong system effectiveness. Compared to the state-of-the-art baseline, TowerInstruct-Mistral-7B, an overwhelming win rate (106 vs. 32) can be observed. Summaries of the annotation results for other directions are in Appendix D.3. We also release all detailed human annotations to the community [^4] Detailed human annotations are available at https://huggingface.co/datasets/Calibration-Translation/Calibration-translation-human-eval..
6.3 Sensitivity to Sampling Size
(a) KIWI-XL scores varying sampling size
(b) COMET scores varying sampling size
Figure 5: Average performance variation on the WMT24 dataset across different sampling sizes during training, measured by (a) the reference-free metric KIWI-XL and (b) the reference-based metric COMET. Results are reported for TowerInstruct-Mistral-7B, using supervised fine-tuning, CPO, and our calibration method.
As §3 mentions, we approximately characterize the decoding space through sampling. Therefore, increasing the sampling size for each source sentence is expected to improve effectiveness. To validate this, we conduct experiments on TowerInstruct-Mistral-7B, varying the sampling size during training. Also, we compared our approach with CPO and selected the highest- and lowest-scoring hypotheses as the accepted and rejected hypotheses, respectively.
Figure 5 provides the results measured by both reference-free and reference-based metrics, i.e., CometKiwi-XL and COMET, in different settings when varying the sampling size for each source sentence. It clearly shows that: (1) SFT with the best-scoring hypothesis has limited impact on translation performances; (2) CPO yields clear improvements over supervised fine-tuning, but these improvements plateau as the sample size exceeds 8; (3) our approach manifests substantial improvements over both CPO and SFT. Moreover, as the sampling size increases, continued improvements can be observed, demonstrating greater potential for further advancements. Also, the results support the underlying hypothesis that more accurate modeling of the decoding space facilitates better calibration, which in turn yields greater performance improvements.
7 Conclusion
This paper focuses on the well-known miscalibration problem in translation, where the training and decoding objectives possibly misalign with real-world translation quality. To mitigate this issue, we introduce a simple yet highly effective method to improve the likelihood-quality calibration during training time by directly optimizing the Pearson correlation objective. Extensive experiments demonstrate several key advantages of our method:
Very clear and consistent performance improvements.
Even when applied to state-of-the-art systems such as Tower, our calibration method yields substantial improvements across a wide range of automatic metrics and human evaluations. For example, a calibrated Tower-7B model with beam search achieves state-of-the-art translation performance comparable to — if not exceeding — that of the much larger Tower-70B-v2 model (which employs test-time optimizations such as MBR decoding), while the calibrated Tower-7B model requires substantially less computation.
A unified view of quality optimization and estimation in translation.
Our approach adopts a shared objective for both quality optimization and quality estimation in translation, offering a unified perspective on these two tasks. Our empirical analysis also supports this perspective, indicating a clear correlation between calibration level and translation performance. Experimentally, the calibrated method achieves both state-of-the-art translation performance and accurate quality estimation within a single model, with the latter even surpassing widely used QE models such as CometKiwi — all without relying on human-annotated data.
Enhanced realization for maximum a posterior decoding.
By calibrating the training objective with real-world quality, the effectiveness of widely used efficient MAP decoding rules, such as beam search, is improved. This paper shows that the effectiveness of beam search originally fell far short of Best-of-N sampling in translation quality for the baseline system. However, our method not only substantially improves performance, but also enables beam search to achieve comparable performance to Best-of-N sampling with 100 hypotheses using only a beam size of 5 — dramatically reducing inference cost without sacrificing quality, showing efficiency potential for real-world deployment.
Appendix A Detailed Dataset Description
WMT22 and WMT24 translation datasets.
In this paper, we use the WMT22 and WMT24 datasets in the general translation track. WMT24 covers 11 language directions. In this paper, we focus on the 9 out-of-English translation directions (ende|es|ru|zh|fr|nl|it|pt|ko) and use them to conduct experiments on Tower series models. All testsets in WMT24 are paragraph-level and share the same English parts, consisting of 960 samples for each direction. WMT22 consists of 22 language directions, covering out-, into-, and non-English translations. In this paper, for a fair comparison, we focus on 10 directions that are covered by ALMA series models, i.e., both out-of-English (ende|cs|is|zh|ru) and into-English(ende|cs|is|zh|ru) translations. Each direction contains 2037 sentence pairs.
WMT22 quality estimation dataset.
To evaluate the effectiveness of using our model’s likelihood as a quality estimation metric, we use the WMT 2022 Quality Estimation (QE) dataset, which provides source sentences, machine-translated outputs, and corresponding human annotations at both the sentence and word levels. The dataset includes direct assessment scores (DA), post-editing effort indicators such as HTER, word-level quality tags (OK/BAD), and multidimensional quality annotation (MQM). In this paper, we use MQM scores as the ground-truth human annotations against which the metrics’ scores are evaluated, as MQM scores come from expert translators and are more reliable than the crowdsourced DA scores. WMT22 QE dataset covers three language directions, i.e., ende, enru, zhen. We focus on the first two as they overlap with those during the Tower base model pretraining, amounting to approximately 1,000 segments per language pair. More detailed dataset descriptions can be found in  [Zerva et al., 2022].
Appendix B More Results for Section 5.1
In this appendix, we provide additional results from the experiments presented in the main text, including (1) a broader range of translation evaluation metrics, (2) the experimental results based on the ALMA model, and (3) the experimental results under on-policy training dynamics.
B.1 Off-policy Results based on Tower in KIWI-XXL and COMET
Table 3: Evaluation of enxx translation on WMT24 using CometKiwi-XXL and COMET. Results are reported for all languages covered during Tower-v1 pretraining. Note that the Tower-v2 models, including Tower-70B-v2, have not been publicly released. For GPT-4o and GPT-4o-mini, we use the prompts following  Hendy et al. [2023]. Notably, according to Kocmi et al. [2024b], improvements of $\geq 1.99$ in XCOMET or $\geq 0.94$ in COMET scores correspond to at least 90% estimated accuracy in human judgment — both of which are achieved by our method.
ende enes enru enzh enfr
Models KIWI-XXL COMET KIWI-XXL COMET KIWI-XXL COMET KIWI-XXL COMET KIWI-XXL COMET
Closed GPT-4o-mini 76.4 82.7 76.3 83.8 75.5 82.5 75.8 84.6 74.7 81.5
GPT-4o 77.7 82.5 77.3 83.8 77.6 82.8 77.6 84.5 76.2 81.7
Tower-70B-v2
Tower-70B-v2 + MBR/TRR
TowerInstruct-7B 76.5 81.2 76.3 82.8 75.9 81.1 74.8 83.1 76.7 81.2
TowerBase-7B
     + SFT on BoN data 77.2 81.3 75.8 82.4 76.2 80.9 74.4 82.4 76.2 81.0
     + CPO 78.9 82.2 78.0 83.2 78.8 82.2 77.8 83.4 78.7 81.2
     + Calibration 79.5 82.8 79.8 83.7 80.4 82.9 78.0 83.2 80.2 81.7
[…]
TowerBase-13B
     + SFT on BoN data 79.0 82.3 77.0 83.1 78.4 82.0 76.8 83.8 77.2 81.5
     + CPO 79.1 82.1 78.6 82.5 80.3 82.6 78.0 83.4 79.3 81.5
     + Calibration 81.3 83.4 80.9 84.1 82.3 83.8 80.4 84.5 81.5 82.2
[…]
     + SFT on BoN data 78.3 82.0 77.5 82.9 78.3 81.5 77.3 84.0 77.3 81.4
     + CPO 79.6 82.2 79.9 83.3 80.5 82.7 79.7 84.8 79.9 81.8
     + Calibration 80.7 83.1 80.6 83.9 82.0 83.6 80.4 84.9 80.8 82.1
ennl enit enpt enko Avg.
Models KIWI-XXL COMET KIWI-XXL COMET KIWI-XXL COMET KIWI-XXL COMET KIWI-XXL COMET
Closed GPT-4o-mini 78.3 84.6 74.1 83.6 77.9 81.9 81.2 86.2 76.7 83.5
GPT-4o 80.7 84.6 76.0 83.8 79.1 81.9 82.3 86.2 78.3 83.5
Tower-70B-v2
Tower-70B-v2 + MBR/TRR
TowerInstruct-7B 81.1 84.4 77.7 83.7 77.9 81.8 80.0 84.7 77.4 82.7
TowerBase-7B
     + SFT on BoN data 80.5 83.5 76.9 83.4 78.6 81.5 82.3 85.3 77.6 82.4
     + CPO 81.9 83.8 79.4 83.7 80.5 81.8 83.7 85.8 79.7 83.0
     + Calibration 83.6 84.8 81.0 84.3 81.9 82.7 84.6 86.1 81.0 83.6
[…]
TowerBase-13B
     + SFT on BoN data 80.8 84.3 77.9 83.8 79.5 81.7 83.6 85.7 78.9 83.1
     + CPO 82.5 84.2 80.2 83.8 79.2 80.7 85.0 86.5 80.2 83.0
     + Calibration 84.5 85.1 82.1 84.6 82.8 82.7 86.2 87.1 82.4 84.2
[…]
     + SFT on BoN data 81.4 84.2 78.4 83.7 79.6 81.7 83.8 86.1 79.1 83.0
     + CPO 83.9 84.8 80.7 84.0 81.9 82.2 85.9 86.9 81.3 83.6
     + Calibration 84.6 85.2 82.3 84.8 83.2 83.0 86.9 87.3 82.4 84.2
Table 3 shows off-policy results measured by two other metrics, i.e., CometKiwi-XXL (abbreviated as KIWI-XXL) and COMET-22 (abbreviated as COMET). Very strong average performance improvements can be observed. For instance, +3.6 and +1.1 points of KIWI-XXL and COMET average gains are shown over TowerInstruct-Mistral-7B.
B.2 Off-policy Results based on ALMA
Table 4: Out-of-English translation evaluation results for ALMA models on the WMT22 dataset, measured by 3 metrics, i.e., COMET, KIWI-XXL, and XCOMET. When applying our calibration method on ALMA-base, consistent substantial improvements across all metrics and directions can be observed.
ende encs COMET KIWI-XXL XCOMET COMET KIWI-XXL XCOMET ALMA-7B-LoRA 85.45 80.70 96.49 89.05 82.06 90.82 + SFT on preferred data 85.42 80.44 96.26 89.11 81.28 90.26 + DPO 85.19 80.02 96.22 88.78 81.03 90.12 + CPO (ALMA-13B-R) 86.06 82.77 97.11 89.61 84.81 91.91 + Calibration (ours) 86.22 83.83 97.25 90.24 86.25 93.06 enis enzh COMET KIWI-XXL XCOMET COMET KIWI-XXL XCOMET ALMA-7B-LoRA 85.44 81.51 89.94 84.87 77.14 88.11 + SFT on preferred data 85.19 80.25 89.15 85.36 78.16 88.34 + DPO 85.20 80.42 88.97 84.73 76.96 87.72 + CPO (ALMA-13B-R) 85.80 82.35 89.63 85.89 81.79 89.55 + Calibration (ours) 86.29 84.11 91.62 86.47 83.18 90.75 enru Avg. COMET KIWI-XXL XCOMET COMET KIWI-XXL XCOMET ALMA-7B-LoRA 87.05 82.60 92.98 86.37 80.80 91.67 + SFT on preferred data 86.88 81.79 92.57 86.39 80.38 91.32 + DPO 86.70 81.61 92.53 86.12 80.01 91.11 + CPO (ALMA-13B-R) 87.86 84.97 94.15 87.04 83.34 92.47 + Calibration (ours) 88.41 86.30 94.46 87.53 84.73 94.43
Table 5: Into-English translation evaluation results for ALMA models on the WMT22 dataset, measured by 3 metrics, i.e., COMET, KIWI-XXL, and XCOMET. When applying our calibration method on ALMA-base, consistent substantial improvements across all metrics and directions can be observed.
deen csen COMET KIWI-XXL XCOMET COMET KIWI-XXL XCOMET ALMA-7B-LoRA 83.95 82.58 92.35 85.93 81.42 81.34 + SFT on preferred data 84.39 82.72 93.19 86.17 81.95 84.55 + DPO 84.02 82.47 92.26 85.87 81.30 81.10 + CPO (ALMA-7B-R) 84.61 83.11 93.85 86.29 82.29 85.76 + Calibration (ours) 85.03 84.30 94.33 86.52 84.13 84.79 isen zhen COMET KIWI-XXL XCOMET COMET KIWI-XXL XCOMET ALMA-7B-LoRA 86.09 84.65 75.02 79.78 73.65 83.94 + SFT on preferred data 86.47 85.23 78.87 80.50 74.91 89.81 + DPO 85.96 84.44 75.19 79.91 73.51 89.22 + CPO (ALMA-7B-R) 86.66 85.13 79.14 80.95 75.72 90.74 + Calibration (ours) 87.29 86.37 79.07 81.33 76.88 92.61 ruen Avg. COMET KIWI-XXL XCOMET COMET-22 KIWI-XXL XCOMET ALMA-7B-LoRA 84.84 80.19 88.50 84.12 80.50 84.23 + SFT on preferred data 85.00 80.47 89.54 84.51 81.06 87.19 + DPO 84.71 80.04 88.34 84.09 80.35 85.22 + CPO (ALMA-7B-R) 85.11 80.69 90.10 84.72 81.39 87.92 + Calibration (ours) 85.12 81.87 90.56 85.06 82.71 88.27
Tables 4 and 5 present the experimental results based on the ALMA model under off-policy settings. Consistently substantial improvements across metrics — COMET-22 (abbreviated as COMET), KIWI-XXL, and XCOMET-XXL (abbreviated as XCOMET) — are observed when applying our calibration method. The resulting performance surpasses all baselines across nearly all language directions, including the strongest ALMA variant, ALMA-7B-R. For example, we observe gains of +0.5 COMET, +1.4 KIWI-XXL, and +2.0 XCOMET points on out-of-English translation compared to ALMA-7B-R.
B.3 On-policy Results based on Tower
Table 6: On-policy results of enxx translation on WMT24 using CometKiwi-XL and XCOMET. Results are reported for all languages covered during Tower-v1 pretraining.
ende enes enru enzh enfr
Models KIWI-XL XCOMET KIWI-XL XCOMET KIWI-XL XCOMET KIWI-XL XCOMET KIWI-XL XCOMET
Closed GPT-4o-mini 68.3 91.7 70.2 87.0 68.1 81.6 69.0 79.7 65.6 83.0
GPT-4o 68.6 92.6 70.6 87.7 69.1 83.4 69.9 81.3 66.0 83.9
Tower-70B-v2
Tower-70B-v2 + MBR/TRR 72.3 74.5 74.2 72.6
TowerInstruct-7B 69.0 91.7 70.8 86.9 69.0 81.5 68.5 78.7 67.9 84.1
TowerBase-7B
     + SFT on BoN data 69.8 92.0 71.6 87.2 69.2 81.8 68.6 78.6 68.7 84.0
     + SFT on CPO 70.2 92.1 72.5 89.0 71.8 84.6 70.0 80.4 69.4 86.3
     + Calibration 71.6 93.6 74.0 90.1 73.1 85.7 71.6 81.8 70.9 87.5
[…]
TowerBase-13B
     + SFT on BoN data 70.6 92.9 71.9 88.8 71.2 84.3 70.3 81.6 68.3 86.0
     + CPO 70.5 92.4 73.2 88.4 72.0 84.4 70.6 82.1 68.8 85.9
     + Calibration 71.5 94.2 74.3 90.3 73.3 86.2 72.0 83.9 70.2 87.3
[…]
     + SFT on BoN data 70.4 92.8 72.5 88.4 71.0 84.3 70.7 81.2 69.0 85.6
     + CPO 70.7 92.4 72.4 88.0 72.1 84.2 71.7 82.5 69.3 85.6
     + Calibration 71.8 93.9 74.1 90.0 73.5 86.3 72.5 84.0 70.5 87.0
ennl enit enpt enko Avg.
Models KIWI-XL XCOMET KIWI-XL XCOMET KIWI-XL XCOMET KIWI-XL XCOMET KIWI-XL XCOMET
Closed GPT-4o-mini 69.4 88.9 68.1 83.7 71.2 87.6 73.2 84.2 69.2 85.3
GPT-4o 70.6 90.5 68.7 85.7 71.5 88.5 73.7 85.6 69.8 86.6
Tower-70B-v2
Tower-70B-v2 + MBR/TRR
TowerInstruct-7B 71.5 90.9 71.1 86.1 71.1 86.8 73.6 82.8 70.3 85.5
TowerBase-7B
     + SFT on BoN data 71.8 90.9 71.5 86.3 71.2 87.0 74.5 82.8 70.8 85.6
     + CPO 72.2 91.2 72.5 87.7 73.7 88.9 75.3 86.4 71.8 87.4
     + Calibration 73.7 92.3 73.9 88.7 75.0 90.5 76.7 87.5 73.4 88.6
[…]
TowerBase-13B
     + SFT on BoN data 72.0 91.6 71.2 87.8 72.6 88.6 75.6 86.8 71.5 87.6
     + CPO 72.1 91.6 72.7 87.9 73.4 88.6 76.4 87.5 72.2 87.6
     + Calibration 73.5 93.2 74.2 89.5 74.4 90.1 77.4 89.4 73.4 89.3
[…]
     + SFT on BoN data 72.5 91.7 72.3 87.7 72.8 88.6 75.5 86.9 71.8 87.5
     + CPO 73.4 92.3 73.0 87.8 72.9 88.3 77.4 88.3 72.5 87.7
     + Calibration 73.9 93.2 74.0 89.3 74.7 90.2 77.7 89.4 73.6 89.3
Table 7: On-policy results of enxx translation on WMT24 using CometKiwi-XXL and COMET. Results are reported for all languages covered during Tower-v1 pretraining.
ende enes enru enzh enfr
Models KIWI-XXL COMET KIWI-XXL COMET KIWI-XXL COMET KIWI-XXL COMET KIWI-XXL COMET
Closed GPT-4o-mini 76.4 82.7 76.3 83.8 75.5 82.5 75.8 84.6 74.7 81.5
GPT-4o 77.7 82.5 77.3 83.8 77.6 82.8 77.6 84.5 76.2 81.7
Tower-70B-v2
Tower-70B-v2 + MBR/TRR
TowerInstruct-7B 76.5 81.2 76.3 82.8 75.9 81.1 74.8 83.1 76.7 81.2
TowerBase-7B
     + SFT on BoN data 76.7 81.4 76.6 82.4 76.1 81.7 75.4 82.2 77.3 81.2
     + CPO 78.3 81.1 79.7 82.7 80.0 82.1 77.8 82.9 79.8 80.9
     + Calibration 79.8 82.2 80.8 83.4 81.2 83.2 79.4 83.8 81.4 81.8
[…]
TowerBase-13B
     + SFT on BoN data 78.7 81.5 78.5 83.4 79.6 82.2 78.4 83.4 78.7 81.6
     + CPO 79.4 82.7 79.7 82.5 80.9 82.8 79.2 83.2 79.7 81.0
     + Calibration 81.0 83.2 81.4 83.4 82.4 83.5 80.7 84.3 81.3 81.7
[…]
     + SFT on BoN data 78.6 82.5 78.6 83.1 79.0 82.4 77.9 84.3 78.6 81.9
     + CPO 78.7 81.9 79.1 82.2 80.3 82.5 79.4 84.3 79.7 81.3
     + Calibration 80.7 83.0 81.3 83.6 82.0 83.6 80.6 84.6 81.2 81.8
ennl enit enpt enko Avg.
Models KIWI-XXL COMET KIWI-XXL COMET KIWI-XXL COMET KIWI-XXL COMET KIWI-XXL COMET
Closed GPT-4o-mini 78.3 84.6 74.1 83.6 77.9 81.9 81.2 86.2 76.7 83.5
GPT-4o 80.7 84.6 76.0 83.8 79.1 81.9 82.3 86.2 78.3 83.5
Tower-70B-v2
Tower-70B-v2 + MBR/TRR
TowerInstruct-7B 81.1 84.4 77.7 83.7 77.9 81.8 80.0 84.7 77.4 82.7
TowerBase-7B
     + SFT on BoN data 81.4 84.9 78.0 84.5 78.7 82.5 80.2 85.0 77.8 82.9
     + CPO 82.6 83.9 80.9 83.8 81.2 82.3 83.3 85.5 80.4 82.8
     + Calibration 84.1 84.7 82.1 84.9 82.9 83.5 84.4 86.2 81.8 83.7
[…]
TowerBase-13B
     + SFT on BoN data 82.2 84.4 80.2 85.1 80.4 83.3 83.9 85.8 80.1 83.4
     + CPO 83.2 83.7 80.5 84.1 81.4 82.3 85.1 86.5 81.0 83.2
     + Calibration 84.8 84.6 82.2 84.6 83.0 83.0 86.5 87.1 82.6 83.9
[…]
     + SFT on BoN data 82.2 84.7 79.8 84.6 80.2 82.6 83.3 86.3 79.8 83.6
     + CPO 83.6 84.6 80.8 84.0 80.8 82.3 85.6 86.7 80.9 83.3
     + Calibration 84.8 84.9 82.1 84.7 82.7 83.1 86.1 87.2 82.4 84.1
Table 6 and Table 7 present on-policy results measured by KIWI-XL and COMET, and by KIWI-XXL and XCOMET, respectively. We found that in the on-policy setting, our calibration method performs better with a relatively smaller learning rate (1e-5), whereas in the off-policy setting, a larger learning rate (5e-5) yields the best performance. All other training settings are the same as those for off-policy, except for sampling from the model itself instead of an external model like GPT-4o-mini.
Appendix C More Results for Section 5.2
[…]
In this appendix, we provide additional results for Section C presented in the main text, including (1) another evaluation direction, namely, enru, and (2) another statistic, Kendall’s $\tau$ for metric meta-evaluation.
Figure 6 shows the Spearman coefficient and the corresponding translation performance in the enru direction. Meanwhile, Figures 7 and 8 present the results using Kendall’s $\tau$ for the ende and enru directions, respectively. It is clear that the main findings, as mentioned in Section 5.2, hold across language directions and statistics.
Appendix D Human Study
In this appendix, we present the detailed human evaluation results for three additional target languages: German, Russian, and Dutch. Section D.1 outlines the annotation criteria used, and Section D.3 reports the score distributions and aggregated summaries for these language directions.
D.1 Annotation Criteria
We follow the standard of the official assessment system during WMT22 [Kocmi et al., 2022], Annotators are asked to rate a pair of source sentences and the corresponding hypothesis with scores ranging from 0 to 6. The descriptions for each quality level are as follows:
  • 0: Nonsense/No meaning preserved: Nearly all information is lost between the translation and the source. Grammar is irrelevant.
  • 2: Some meaning preserved: The translation preserves some of the meaning of the source but misses significant parts. The narrative is hard to follow due to fundamental errors. Grammar may be poor.
  • 4: Most meaning preserved and few grammar mistakes: The translation retains most of the meaning of the source. It may have some grammar mistakes or minor contextual inconsistencies.
  • 6: Perfect meaning and grammar: The meaning of the translation is completely consistent with the source and the surrounding context (if applicable). The grammar is also correct.
Note that for WMT, a continuous score is allowed for annotators. However, we require the annotator to score with integers. For example, a score of 5 means in between of minor error and perfection, which may be reflected in tone, grammar, or some subtle wording.
D.2 English-to-Chinese Translation Annotation Analysis
Figure 9: Detailed annotation results illustrate the number of changes in enzh translations. Each data point represents the number of samples corresponding to a specific pair of scores before and after calibration (vertical and horizontal axes), respectively.
Figure 9 presents detailed annotation results, illustrating the changes in enzh translations before (TowerInstruct-Mistral-7B) and after applying our calibration method. We can see that:
  • Both systems perform well, as only a few translations received scores below 3.
  • The primary improvements over the baseline are observed in better handling of minor to major errors. For instance, 29 and 47 translations previously rated 4 and 5 (minor errors) were upgraded to 6 (nearly perfect), respectively.
Note that annotators were blinded to the source of each translation to prevent any systematic bias.
D.3 Other Annotation Results
Annotation results in other translation directions are in preparation and will be published soon. Also, we will release all detailed human annotations to the community [^5] Detailed human annotations are available at https://huggingface.co/datasets/Calibration-Translation/Calibration-translation-human-eval..
Appendix E Prompt Templates
The prompts used in our experiments are presented in Sections E.1, E.2, and E.3. To ensure a fair comparison, the prompts strictly follow the default settings of the original models [Xu et al., 2023, Rei et al., 2024a] and prior work [Xu et al., 2024].
E.1 GPT Prompt
System: You are a helpful translator and only output the result. User: ### Translate this sentence from <source language> to <target language>, <source language>: <source sentence> ### <target language>:
E.2 Tower Prompt
Translate the following text from <source language> into <target language>. <source language>: <source sentence> <target language>:
E.3 ALMA Prompt
Translate this from <source language> into <target language>: <source language>: <source sentence> <target language>: